ThatCreepyBoy:

For my brother who's too lazy to get on. Find the sum of a finite geometric sequence from n = 1 to n = 6, using the expression −2(5)n − 1. 1,223 −1,023 7,812 −7,812

Vocaloid:

take the expression −2(5)n − 1 plug in n = 1, n = 2, n = 3... until n = 6 then add up the results

ThatCreepyBoy:

So it would be -2(5)1 - 1? :O if I were to replace n with 1?

Vocaloid:

yes

ThatCreepyBoy:

I got -10

ThatCreepyBoy:

-2?

Allison:

I got D. -7,812

moreso:

if its a geometric sequence shouldn't it be $$ -2(5)^{n-1}$$

Allison:

Am I right @Vocaloid ?

Vocaloid:

oof morseo I think you're right

ThatCreepyBoy:

-2 * 5 = -10, -10 * 1 -10?

Vocaloid:

-2(5)^(1-1) = ?

ThatCreepyBoy:

Sorry :P

moreso:

We want \[\sum_{n=1}^{6} -2(5)^{n-1} = -2(5)^{1-1} + -2(5)^{2-1} +-2(5)^{3-1} \\ + -2(5)^{4-1} +-2(5)^{5-1} +-2(5)^{6-1}\]

moreso:

Can you add those?

ThatCreepyBoy:

I'll have to wait till I get home.

Allison:

It's D e.e